# Optimization under Attack: Resilience, Vulnerability, and the Path to Collapse

### v1

@@ V-B Pareto Optimality Analysis
Pareto optimality analysis identifies Pareto fronts knee points that represent optimal system performance under varying adversarial conditions. Figure 2 highlights system resilience to adversarial severity and proportions across the Energy, Voting, and Privacy datasets. Figure 2(a) focuses on the impact of adversarial severity ($\beta$) on tolerated scale of adversaries, while Figure 2(b) explores how varying adversary densities affect tolerated behavior levels.
[…]
The Voting dataset demonstrates robust stability. Figure 2(a) shows consistent knee points tolerating 50–60% adversary agents across a broad range of behavior levels ($0.1\leq\beta\leq 0.7$). Even under high adversarial severity ($0.86\leq\beta\leq 0.9$), as illustrated in Figure 2(b), the system exhibits exceptional resilience, tolerating densities from 10% to 100%. This highlights the dataset high tolerance to adversarial influence.
The Privacy dataset exhibits slightly difference dynamics between high and low privacy-preserving signals. In Figure 2(a), tolerated scale of adversaries for high signal configurations are approximately 70% at $\beta<0.5$, but drop to 50% at higher intensities. The low signal Privacy scenario displays smoother declines, with tolerated densities averaging 68% Adversary agents across behavior levels. Figure 2(b) shows that high signal configurations begin at knee points of $\beta=0.03$ for 10% adversary agents, rising gradually to $\beta=0.2$ at 50–90%. Low privacy-preserving signal exhibits a smoother gradient, peaking at $\beta=0.36$ at 20% Adversary agents and declining to $\beta=0.13$ at full adversarial saturation.
@@ V-C Structure Analysis
This section analyzes how hierarchical structures influence inefficiency costs in Energy, Voting, and Privacy datasets. This analysis is divided into two approaches: individual layer-wise and cumulative structure analysis.
@@ V-C1 Layer-wise structural analysis
Figure 3 highlights inefficiency costs across hierarchical layers for varying scale of adversaries at 25%, 50%, 75%, and 100% and behavior levels ($\beta$) at 0.1, 0.4, 0.7, and 1.0. In all datasets, deeper layers and layers with higher agent proportions exhibit greater inefficiency escalation under intensified adversarial influence.
In the Energy dataset (Figure 3(a)), with 1,000 agents distributed across 10 layers, inefficiency costs remain stable up to Layer 5 at moderate adversarial severity ($\beta=0.1$, 0.4, and 0.7). Beyond Layer 6, costs fluctuate and peak at Layers 7 and 8. Notably, while Layer 9 contains the highest number of agents (256), inefficiency only peaks at $\beta=1.0$ and 100% Adversary agents, highlighting the compounded vulnerability of deeper layers under extreme adversarial influence. This emphasizes the increasing impact of both adversary density and adversarial severity on deeper layers.
The Voting dataset (Figure 3(b)), with 266 agents distributed across 9 layers, shows minimal inefficiency in the upper layers (1–5) across all adversarial configurations. However, inefficiency rises sharply beyond Layer 5, with Layer 7 consistently incurring the highest costs, especially at higher adversarial severity ($\beta\geq 0.7$) and scale of adversaries of 75% and 100%. At $\beta=1.0$, inefficiency increases steadily from the root (Layer 1) to the leaves (Layer 9), highlighting the compounding impact of adversarial behavior as it propagates deeper into the hierarchy.
The Privacy dataset (Figures 3(c) and 3(d)), with 72 agents distributed across 7 layers, exhibits almost similar trends for high and low privacy-preserving signal. At low adversarial behavior ($\beta=0.1$), inefficiency costs remain stable across all layers, demonstrating resilience to mild adversarial influence. At higher levels of adversarial behavior ($\beta\geq 0.7$), inefficiency peaks at Layer 6, which has the highest agent proportion (32 agents). Interestingly, Layer 7, with fewer agents, incurs lower costs even under severe adversarial conditions. Minor deviations, such as Layer 5 occasionally surpassing Layer 6 in the low signal scenario, highlight the nuanced relationship between layer depth, agent density, and adversarial behavior.
These results suggest that inefficiency costs in hierarchical systems are highly sensitive to layer depth, scale of adversaries, and the level of adversarial behavior. Deeper layers and those with higher agent proportion emerge as critical vulnerabilities under intensified adversarial configurations. Reinforcing deeper layers and redistributing agent densities could mitigate inefficiency spikes, especially in applications with high adversarial risks.
(a) Top-down cumulative
(b) Bottom-up cumulative
@@ V-C2 Cumulative structural analysis
Another approach assesses the cumulative impact of adversary agents on inefficiency costs as they propagate through the hierarchical network structure in two configurations: top-down (root-to-leaf) and bottom-up (leaf-to-root). Figure 4 presents inefficiency trends for scale of adversaries and adversarial severity (0.1, 0.4, 0.7, 1.0) across Energy, Voting, and Privacy datasets.
The cumulative structural analysis of the Energy dataset reveals distinct differences in inefficiency trends between top-down and bottom-up propagation. At $\beta=1$, top-down propagation results in significantly higher initial inefficiency costs, reaching 272 with a single adversarial agent, compared to just 0.8 in bottom-up propagation. Inefficiency costs in the top-down configuration also rise more rapidly, indicating accelerated system deterioration. In contrast, bottom-up propagation exhibits a slower emergence of inefficiency, with costs remaining low up to 20% adversarial agent participation, followed by exponential growth. These trends emphasize the critical impact of adversarial agent placement, as costs escalate faster and peak higher when adversaries are positioned in top layers. At lower adversarial severity (0.1, 0.4, 0.7), inefficiency costs remain low and stable, fluctuating between 0.1 and 0.5, even as number of adversaries increases. This highlights the system resilience to moderate adversarial behavior.
The Voting dataset shows distinct trends in inefficiency costs. At $\beta=1$, inefficiency costs start at approximately 2.8, similar to lower $\beta$ levels, but rise sharply as the proportion of adversary agents increases, peaking at 6.0 when adversary agents saturate the network (100%). In the top-down propagation, inefficiency costs accelerate rapidly within the top four layers, driven by early adversarial influence. This is followed by a clear decrease between Layers 5 and 6, corresponding to scale of adversaries ranging from 6% to 17%, indicating that inefficiencies stabilize as the system progresses deeper into the hierarchy. At lower adversarial severity ($\beta=0.1,0.4,0.7$), inefficiency costs remain low and stable, fluctuating slightly between 2.8 and 2.9, regardless of scale of adversaries. Notably, at $\beta=0.7$, inefficiency costs increase faster in the top-down propagation compared to the bottom-up propagation, underscoring the impact of adversary agents propagation direction on system performance under moderate adversarial behavior.
The Privacy dataset reveals resilience at low adversarial severity ($\beta=0.1$), with minimal inefficiency increases (0.01–0.05) across all layers in both top-down and bottom-up propagations, even under full adversary participation. At higher adversarial levels ($\beta=0.7,0.4$ and $\beta=1$), inefficiency costs rise significantly across layers. For high privacy-preserving signal, inefficiency emerges early with just one adversary agent in top-down propagation at Layer 1 with $\beta=0.4,0.7$, and $1$, while in bottom-up propagation, inefficiency begins with low value $<62.8$. In both propagations Layer 2 shows heightened sensitivity, where even 2-3 adversary agents sharply increase inefficiency. Layer 6 experiences notable surges when the scale of adversaries exceeds 50% for both propagation directions. In low privacy-preserving signals, inefficiency costs rise more prominently in top-down propagation, particularly at Layer 5, with sharp increases at $\beta=0.4,0.7$, and $1$ when scale of adversaries exceeds 25%. In bottom-up propagation, inefficiency escalates at $\beta=0.7$ and $1$ when scale of adversaries exceeds 70%. These patterns emphasize structural vulnerabilities, with Layers 2, 5, and 6 emerging as critical points for inefficiency escalation under intensified adversarial influence.
V-D Discussion
The results of this study reveal critical insights into adversarial influence on multi-objective distributed optimization systems, uncovering opportunities for designing self-healing strategies that enhance system resilience. By examining the dynamics of resilience, vulnerability, and collapse across diverse datasets, the analysis offers valuable lessons for addressing challenges in adversarial environments.
The Energy and Voting datasets exhibit strong resilience, withstanding a broad range of adversarial configurations and transitioning smoothly from vulnerability to collapse under higher adversarial severity. In contrast, the Privacy dataset demonstrates heightened sensitivity to adversarial scales and severity, particularly under high privacy-preservation constraints. This results underscore the importance of tailoring optimization strategies, particularly for privacy-preserving systems.
Pareto analysis identifies critical resilience thresholds where systems can balance inefficiency and discomfort costs. The Energy dataset sustains resilience up to 90% adversarial scale, collapsing only under extreme severity, while the Voting dataset maintains consistent performance across configurations, showcasing robustness. The Privacy dataset requires earlier interventions for high privacy-preserving configurations but allows for more gradual adjustments under low privacy-preserving scenarios. These findings reinforce the importance of identifying Pareto knee points to enable dynamic trade-off prioritization and effective system adaptation under real-time adversarial conditions.
Structural analysis reveals vulnerabilities in deeper hierarchical layers with higher number of agents, which are more prone to inefficiency propagation under intensified adversarial influence. Targeted strategies, such as reinforcing critical layers, redistributing agent densities, and employing localized containment mechanisms, can mitigate inefficiencies. Top-down configurations show rapid cascading of adversarial impact from top layers, demanding robust protections at higher levels, whereas bottom-up configurations exhibit slower propagation, suggesting inherent resilience.
Integrating adaptive fault-monitoring mechanisms is essential for maintaining system stability. Dynamically detecting transitions between resilience, vulnerability, and collapse allows for strategic resource allocation. For instance, energy systems can employ load-balancing strategies to address inefficiency, while hierarchical vulnerabilities in distributed optimization algorithms should be addressed through structural reinforcements and proactive monitoring.
This research provides a comprehensive framework for understanding adversarial impacts on distributed optimization systems and offers actionable insights for designing robust, self-healing mechanisms. By combining structural interventions, adaptive monitoring, and informed resource allocation, distributed systems can maintain efficiency and resilience even under significant adversarial configurations.
@@ VI Conclusion and Future work
This study provides a comprehensive analysis of resilience, vulnerability, and collapse dynamics in multi-agent distributed optimization systems under adversarial conditions. By systematically examining the impact of the scale of adversaries, adversarial severity, and hierarchical structure, it identifies critical thresholds where system performance transitions from resilience to collapse. These insights offer practical strategies to enhance system design and resilience against adversarial influences.
The results highlight critical thresholds where systems transition from resilience to collapse, emphasizing the interplay of structural vulnerabilities and adversarial factors. These findings provide a foundation for designing tailored resilience strategies across diverse operational contexts.
A significant contribution of this work is the introduction of a benchmark dataset and software artifact for I-EPOS collective learning. These resources enable researchers to further investigate adversarial dynamics and develop advanced self-healing strategies. Integrating real-time detection and mitigation mechanisms will be critical for dynamically restoring system performance and resilience in the presence of adversarial behaviors.
Future work will focus on embedding real-time adaptive mechanisms within practical systems to monitor and counteract adversarial influences dynamically. Exploring complex network topologies, diverse adversarial scenarios, and broader application domains will deepen understanding and support the development of robust, fault-tolerant distributed optimization systems capable of maintaining efficiency and stability under challenging conditions.

### v2

@@ V Results Analysis and Discussion
V-A Resilience Analysis
Figure 1 presents inefficiency, discomfort, and compromised discomfort costs across varying adversary scales and severity levels ($\beta$). These metrics capture degradation in system performance, agent satisfaction, and the impact on legitimate agents. The graphs include resilience, vulnerability, and collapse thresholds, with overlaid Pareto knee points. For the inefficiency and compromised‐discomfort plots, knees mark the best trade‐off between those two metrics; on the discomfort heatmaps, the knee indicates the minimal total discomfort for a given inefficiency.
Inefficiency cost (Figure 1(a)) remains low in the resilience zone, especially when adversary ratios are below 30–50% and $\beta<0.8$. As adversary scale and severity increase, systems shift from Resilience to Vulnerability and Collapse, with thresholds varying across datasets. Energy and voting maintain  90% resilience and only 3–5% collapse, while privacy configurations show earlier collapse near 20%. In energy, inefficiency peaks at $\beta=1$ near 4000 cost  [^10] Values for $\beta=1$ in the energy dataset are excluded from visualizations to avoid heatmap saturation due to extremely high inefficiency values, when adversaries exceed 85%. In the voting dataset, inefficiency increases by 111% in the vulnerable region, with collapse triggered beyond 70% adversaries at $\beta\geq 0.96$. Privacy collapses earlier: $\beta\geq 0.3$ at 50% adversaries in the high privacy-preserving signal with 50% adversaries, and $\beta\geq 0.5$ in the low privacy-preserving signal.
Discomfort cost (Figure 1(b)) shows a consistent decline as adversarial intensity increases. In resilient regions, typically below 30–50% adversary presence and $\beta<0.7$, discomfort initially remains high but declines gradually, then sharply in collapse regions, as adversary scale and severity increase. In the energy dataset, discomfort remains high only below 30% adversaries and drops by 8% before collapsing to near-zero. The voting dataset follows a similar pattern, with discomfort gradually decreasing and fully eliminated in collapse. The privacy datasets exhibit earlier collapse ($\beta>0.3$, $>$60% adversaries), resulting in faster discomfort decline.
Compromised discomfort cost (Figure 1(c)) increases with adversarial presence. In the energy dataset, it remains low for $\beta<0.8$ even when all agents are adversarial, but rises sharply when the adversary scale exceeds 40% at $\beta>0.9$. In voting, the cost remain low under mild attacks ($\beta<0.2$, adversary scales $<20\%$) and increase gradually beyond 50% at $\beta>0.7$. Privacy datasets show a steeper rise, peaking at 9 (high privacy) and 10 (low privacy) when 90% of agents are adversarial. Resilience disappears at high adversary densities ($>80\%$), even under low severity, and collapse emerges at low $\beta$ when agent compromise is widespread. While energy and voting are more sensitive to $\beta$, privacy is primarily affected by adversary scale.
Interestingly, a noticeable resilience lag exists between the system-level inefficiency and the agent-level discomfort and compromise costs. In all datasets, a substantial portion of configurations classified as ”Vulnerable” or even ”Collapsed” by inefficiency remain in the ”Resilient” state when evaluated by discomfort metrics. For instance, in the energy dataset, while only 6% of configurations are in vulnerability based on inefficiency, 59% fall under vulnerability based on discomfort—indicating that discomfort degrades far earlier. Similarly, in the voting dataset, 21% of configurations are classified as vulnerable by inefficiency, compared to 70% by discomfort. This pattern reveals a lag of up to 40–60% in the transition from individual to system-level degradation, with discomfort and compromise metrics acting as early-warning indicators long before global inefficiency surfaces. This insight underlines the importance of incorporating agent-centric metrics for proactive resilience monitoring.
@@ V-B Pareto Optimality Analysis
Pareto optimality analysis identifies fronts and knee points reflecting optimal trade-offs under varying adversarial conditions. Figure 2 shows how system inefficiency relates to discomfort experienced by legitimate agents across different severity levels and adversary scales.
Figure 2(a) focuses on the impact of adversarial severity ($\beta$) on the tolerated scale of adversaries. The voting dataset maintains stable knee points across 50–60% adversary ratios over a broad severity range ($0.1\leq\beta\leq 0.7$), with consistent fronts even at high $\beta$. The privacy dataset shows distinct patterns: the high privacy-preserving signal tolerates 70% adversaries at $\beta<0.5$, decreasing to 50% at higher severities; the low privacy-preserving signal exhibits smoother transitions with average tolerance near 68%.
Figure 2(b) evaluates the impact of adversarial scale on tolerated severity. For visual clarity, only 20 adversary scales are shown per dataset. In energy, knee points are stable at $0.30.03<\beta<$ 0.3 for up to 90% adversaries, increase to $\beta=0.9$ at lower scales ($<40\%$), and decline to $\beta<0.1$ at full scale. Voting shows higher tolerance, with knee points extending $\beta=0.8$ even under full adversaries. Privacy dataset shows greater variability: the high privacy-preserving configuration reaches $\beta=0.2$ for 50–90% adversaries, while the low privacy-preserving case peaks at $\beta=0.36$ for 20% adversaries and declines to $\beta=0.13$ at full scale. Detailed Pareto plots are in Appendix Appendix B: Pareto Optimality Visualizations.
@@ V-C Structure Analysis
This section analyzes how hierarchical structures influence inefficiency costs across datasets using two approaches: layer-wise and cumulative structural analysis.
@@ V-C1 Layer-wise Structural Analysis
Figure 3 presents the inefficiency costs across hierarchical layers under adversary scales of 25%, 50%, 75%, and 100%. For layers with few agents, intermediate adversary ratios are approximated by averaging to the closest feasible configurations.
In the energy dataset (Figure 3(a)), with 1,000 agents over 10 layers, inefficiency remains low (below 0.16) across moderate adversarial scales and gradually increases with severity ($\beta<1.0$). The minimum inefficiency occurs at layer 2 under 100% adversaries and $\beta=0.56$, while inefficiency peaks at layer 8 under $\beta=0.96$, marking a 144% increase from the minimum. Profiles remain smooth under 25% and 50% adversary scales but fluctuate more at 75% and 100%. Under extreme severity ($\beta=1$), inefficiency surges sharply, reaching over 723 at full adversarial saturation. With this severity, the root layer shows moderate inefficiency, exceeding layers 2–7 at 75% and matching layers 2 and 3 at 100%, followed by a steady increase down the hierarchy. Despite layer 10 hosting the largest agent population, its inefficiency values are slightly lower than those of layer 9.
In the voting dataset (Figure 3(b)), with 266 agents over 9 layers, inefficiency remains stable across all layers under low adversarial severity ($\beta\leq 0.3$). As severity increases, costs escalate, particularly at the root (Layer 1) and Layers 2 and 3, which consistently experience higher inefficiency than deeper layers. The maximum inefficiency is observed at layer 8 under 100% adversaries, where agent density is highest. Layer 9,despite its depth, shows lower inefficiency due to a smaller agent count. At 25% and 50% adversary scales, inefficiency profiles remain relatively smooth; however, at 75% and 100%, cost escalations become pronounced, particularly in upper and middle layers.
In the privacy dataset (Figures 3(c) and 3(d)), with 72 agents across 7 layers, inefficiency remains low and uniform under low severities. In the high privacy-preserving configuration, the root layer consistently incurs higher costs than subsequent layers across all adversary scales, with inefficiency peaking at layer 6 (the layer with the largest number of agents) under $\beta\geq 0.7$. Layer 7, although deeper, has lower costs due to fewer agents. In the low privacy-preserving configuration, similar trends are observed with occasional fluctuations between layers 5 and 6 at higher adversary scales. Overall, inefficiency profiles remain structurally consistent between both privacy settings, though severity levels accelerate cost increases in the high-privacy case.
@@ V-C2 Cumulative Structural Analysis
Cumulative analysis evaluates how adversarial influence propagates through hierarchical structures in two configurations: top-down (root-to-leaf) and bottom-up (leaf-to-root). Figure 4 shows inefficiency across layers under both directions.
In the top-down positioning, the energy dataset remains resilient up to layer 7 with 20% adversaries. Collapse occurs at layer 10 when the adversary ratio exceeds 50% and $\beta>0.9$. In voting, resilience holds through layer 4, with vulnerability at layer 5 and collapse at layers 6–7 under high severity ($\beta>0.9$). In the privacy datasets, both signals remain resilient in the top four layers, with vulnerability emerging at layer 4 for $\beta>0.7$. Collapse appears at layer 6 in the high privacy-preserving signal with over 60% adversaries, and slightly earlier in the low privacy-preserving signal with 50% adversaries.
In the bottom-up direction, the energy dataset shows a narrower vulnerability and collapse region. Vulnerability begins at layer 9 with 50% adversaries, and collapse follows at layer 10 under high severity ($\beta>0.9$). Voting results are similar across both directions, with transitions driven more by adversary ratio than structural depth. Vulnerability emerges at layer 6 with 20% adversaries, and collapse follows at layer 5 with 60%. In privacy, vulnerability appears at layer 6 in both signals; collapse occurs at layer 6 (high privacy-preserving) and layer 5 (low privacy-preserving).
(a) Inefficiency cost
(b) Discomfort cost
Figure 5: Normalized inefficiency and discomfort costs on synthetic Gaussian data across varying numbers of agents, plan options per agent, and adversary ratios.
V-D Scalability Analysis with Synthetic Data
To benchmark scalability and validate trends from real datasets, we use a synthetic Gaussian dataset to systematically vary agent population (10–100) and plan count (2–10) under different adversarial ratios. This controlled setup reveals optimization performance across configurations under adversarial pressure [^11] Results exclude $\beta=1$ to avoid saturation effects and highlight subtler trends..
Figure 5(a) shows normalized inefficiency costs across agent counts, adversary ratios, and plan numbers. Inefficiency increases notably beyond 60% adversarial presence and is most pronounced in systems with smaller agent populations. Configurations with fewer plans exhibit high inefficiency even under moderate adversarial influence, indicating lower tolerance. At 100% adversaries, inefficiency peaks in 10-agent systems. In contrast, larger agent populations and broader plan spaces significantly enhance resilience, suppressing inefficiency across most adversarial scales.
Figure 5(b) presents normalized discomfort costs under the same conditions. Discomfort increases with the number of plans. Systems with two plans show a gradual decline in discomfort as adversary ratios increase, while those with more plans experience sharper reductions. Discomfort variation expands with plan complexity, indicating greater sensitivity. The number of agents has minimal impact, though the 10-agent setting occasionally shows marginally lower discomfort than larger systems.
V-E Summary of Findings and Discussion
The key findings of this work can be summarized as follows:
  1. 1.
    The interplay of adversarial scale and severity determines the resilience, vulnerability and collapse of distributed multi-objective optimization, which is strongly influenced by the optimization scenario.
  2. 2.
    Distributed multi-objective optimization can predominantly remain resilient, even for high adversarial scales or severity.
  3. 3.
    Adversarial attacks trigger high comfort losses by legitimate agents as collective compromises that reduce the likelihood of collapse for higher vulnerability and resilience.
  4. 4.
    Comfort compromises of legitimate agents for preserving system efficiency under adversarial attacks are predominantly required for high adversarial severity.
  5. 5.
    Pareto optimal points for adversarial severity levels and adversarial scales are mainly in the resilience trajectory. However, Pareto optimal points for high adversarial scales can expand to vulnerability and collapse trajectories.
  6. 6.
    High adversarial scales reduce the comfort compromises required by the legitimate agents in the Pareto optimal points for adversarial severity levels that can be tolerated.
  7. 7.
    High adversarial severity levels reduce the system efficiency in the Pareto optimal points for adversarial scales that can be tolerated.
  8. 8.
    Lower hierarchical levels with higher scales of agents within hierarchical structures of distributed multi-objective optimization are more vulnerable to adversarial attacks than top levels with lower scales of agents.
  9. 9.
    A top-down positioning of adversary agents within hierarchical structures of distributed multi-objective optimization is more impactful on system performance: higher vulnerability, likelihood of collapse, and lower resilience.
  10. 10.
    Systems with a small number of agents and low plan diversity are more susceptible to inefficiency increases under adversarial pressure, even at moderate adversary ratios.
  11. 11.
    Broader plan spaces and larger agent populations enhance system resilience in distributed multi-objective optimization, significantly suppressing inefficiency across adversarial configurations.
The findings of this paper can be used to develop and enhance corrective self-healing strategies [65] that are cost-effective in practice. They can also be used to design incentive mechanisms [66] that ensure agents comply to certain standards of safety in critical infrastructures such as smart grids. For instance, one challenge of fault-correction mechanism is the timely detection of adversary agents to mitigate their impact [65]. Apparently, redundancy mechanisms and rollback operations orchestrated by monitoring mechanisms are resource-intensive [65, 67]. They require frequent checks that involve computations and exchange of messages and they usually rely on static thresholds or even manual operations [67]. This is where the insights of this work can find applicability: these mechanisms can adapt based on the status of the system, for instance, whether it is in the resilience, vulnerability and collapse state. Knowing apriori that an optimization process can tolerate certain adversarial scales and severity can simplify and reduce the costs of applying prevention and mitigation measures. It can also provide new insights for security policies, for instance, prioritizing the protection of agents at the top of the hierarchical optimization structure with stronger security safeguards and resources allocated for this purpose [68]. While creating these strategies does not fall into the scope of this paper, it is part of the future work to pursue.
@@ VI Conclusion and Future work
This study provides a comprehensive analysis of resilience, vulnerability, and collapse dynamics in multi-agent distributed optimization under adversarial conditions. By systematically examining adversary scale, severity, and network structure, we identify critical thresholds where systems transition from stability to failure. These findings offer actionable guidance for designing and enhancing the performance of recovery and healing strategies. A key contribution of this work is the release of a large-scale benchmark dataset, generated from over 112 million experiments using the proposed adversarial model. This dataset supports systematic evaluation of adversarial impacts and facilitates reproducible research across domains of distributed optimization.
Although the adversarial model is designed to be general-purpose, the evaluation can be extended to other algorithms in future work. Additionally, the current experiments model adversarial behavior with static severity levels, which may not fully capture dynamic or strategic adversary actions. Furthermore, resilience has been analyzed under hierarchical network structures, leaving the behavior under alternative topologies an open area for exploration.
Future work will focus on embedding adaptive monitoring and mitigation mechanisms into real-time distributed systems. Investigating more complex network structures, dynamic adversarial strategies, and diverse application domains will further advance the development of robust, fault-tolerant optimization systems capable of maintaining performance under adversarial conditions.